OpenStudy (annie96):

Question 1.1. 2x + 3y + 5x – 8y – 9 (Points : 1) 7x – 5y – 9 7x + 5y – 9 7x - 11 y – 9 7x - 3y - 9 Question 2.2. 6a + (-2c) + (-4a) + c (Points : 1) 10a + 3c 2a + 3c 2a – c 10 + c

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

question 1: 2x+5x=7x 3y-8y=-5y -9 question 2: 6a+(-4a)= 2a -2c+c=-1c

OpenStudy (shadowlegendx):

Remember: X can not be combined with Y Y can not be combined with X Use like terms and combine them correctly to solve the equation.

OpenStudy (annie96):

Thank you guys.

OpenStudy (shadowlegendx):

Brenda did all the work tho xD